On September 25, 2003 an earthquake that
measured 7.4 on the Richter scale shook Hokkaido, Japan.
How much energy (joules) did the earthquake emit?
I get 10^15.5 joules using book formula of M=(2/3)log(E/10^4.4). The answer in my book says 3.16 x 10^15 joules.

Found 2 solutions by Alan3354, math_tutor2020:
Answer by Alan3354(69443)

On September 25, 2003 an earthquake that
measured 7.4 on the Richter scale shook Hokkaido, Japan.
How much energy (joules) did the earthquake emit?
I get 10^15.5 joules using book formula of M=(2/3)log(E/10^4.4). The answer in my book says 3.16 x 10^15 joules.
----------------------
What is the E?
What is the base of the log?

Answer by math_tutor2020(3817)


M = (2/3)*log(E/10^4.4)
7.4 = (2/3)*log(E/10^4.4)
7.4*(3/2) = log(E/10^4.4)
11.1 = log(E/10^4.4)
E/10^4.4 = 10^11.1
E/10^4.4 = 125,892,541,179.417
E = (25,118.8643150959)*(125,892,541,179.417)
E = (2.5118 * 10^4)*(1.2589 * 10^11)
E = (2.5118*1.2589) * (10^4*10^11)
E = 3.16210502 * 10^(4+11)
E = 3.16 * 10^15

Answer: 3.16 * 10^15 joules approximately